OpenStudy (anonymous):

Find the roots without using calculator 6x^4 - 8x^3 + 1 = 0

OpenStudy (anonymous):

everything i can think of involves using a calculator in some way

OpenStudy (anonymous):

there is no negative value and there are two positive answers. 1 close to but a lil over .5 and 1 close to but a lil over 1

OpenStudy (anonymous):

if u were willing to do the math u can use newtons method using .5 and 1 as ur starting points. u wont get an exact answer but im sure u'll get it as far as u need to. Newton's method: \[X(n+1)=X(n)+f(x)/f'(x)\] f(x)=6x^4-8x^3+1 f'(x)=24x^3-24x^2

OpenStudy (anonymous):

without using a calculator you can graph this 6x^4 - 8x^3 + 1 = 0 by graphing you get x=0.614 and x=1.25 ans....

OpenStudy (anonymous):

or try newtons method as jathoday mention...

OpenStudy (anonymous):

i thought graphing itwould b the same as using a calculator. if u can graph it thats the way to go..